# Artificial Liver Classifier: A New Alternative to Conventional Machine Learning Models

### v1

@@ 5 Results
This section presents the performance results of the proposed ALC on multiple benchmark datasets, as described in Section 4.1. The experimental parameter settings were configured for each dataset as follows: 500 detoxification cycles, a detoxification power of 10, and dataset-specific numbers of lobules. Specifically, the number of lobules was set to 10 for Iris Flower and Breast Cancer Wisconsin, 15 for Wine and Voice Gender, and 50 for MNIST. These values were adjusted based on several trials and errors. Moreover, each dataset has been split into a training set comprising 80% of the entire dataset and a validation set comprising 20%. To facilitate later comparison and analysis, additional classifiers, including MLP, SVM, LR, and XGBoost (XGB), were executed on the same datasets. However, all experiments were conducted on an MSI GL63 8RD laptop equipped with an Intel® Core™ i7-8750H × 12 processor and 32 GB of memory. This consistent setup ensured a robust evaluation of the proposed ALC alongside the other classifiers under the same conditions.
@@ 5.1 Performance Metrics
To evaluate the performance of the proposed ALC, several metrics were employed, including log loss (cross-entropy loss), accuracy, precision, recall, F1-score, and training time. Initially, Log loss (Equation 10) quantifies the divergence between predicted probabilities and actual labels, where lower values indicate better predictive performance [60]. The accuracy (Equation 11) measures the proportion of correctly classified instances, serving as a straightforward indicator of overall correctness. Moreover, precision (Equation 12) evaluates the proportion of true positives among all positive predictions, emphasizing the model’s ability to reduce false positives. In contrast, recall (Equation 13) focuses on the proportion of true positives among all actual positive instances, highlighting the importance of minimizing false negatives. Furthermore, the F1-score (Equation 14), as the harmonic mean of precision and recall, provides a balanced assessment when class distributions are imbalanced [61]. Moreover, the overfitting gap defined as the difference between training and validation accuracy, provides insights into generalization. A smaller value indicate better generalization, while a larger value indicates overfitting, where the model excels on the training set but struggles with unseen data. Finally, the training time reflects the duration required to train the model, offering insight into its computational efficiency.
Log $\text{Log Loss}=-\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right)$ (10)
@@
[…]
@@
5.2 Experimental results
The performance results of the proposed ALC are presents through this subsection, summarized in the figures and tables. Additionally, comparisons with other classifiers, including MLP, SVM, LR, and XGB, have been conducted on the five datasets described in Section 4.1.
(a) Log loss on the Iris dataset
(b) Accuracy on the Iris dataset
Table 1: Performance results of the proposed ALC and other classifiers on the Iris Flower dataset.
Classifier ALC XGB SVM MLP LR
Training set results
Loss 0.0514 0.0151 0.0668 0.3129 0.1034
Accuracy 0.9917 1.0000 0.9750 0.9000 0.9750
Precision 0.9919 1.0000 0.9767 0.9057 0.9768
Recall 0.9917 1.0000 0.9750 0.9000 0.9750
F1-Score 0.9917 1.0000 0.9750 0.8996 0.9750
Validation set results
Loss 0.0176 0.0087 0.0704 0.2543 0.0539
@@ 5.2 Experimental results
Overfitting -0.0241% -0.0121% -0.0342% -0.0714% -0.0438%
Time (sec.) 2.18 0.96 4.13 4.28 4.33
Table 1 presents the performance results of the proposed ALC and other classifiers on the Iris Flower dataset, for both the training and validation set. Additionally, Figures 5a and 5b display the loss and accuracy, respectively, on the validation set. The proposed ALC achieved 100% accuracy on the validation set, with a loss of 0.0176 and an overfitting gap of -0.0241%, with a training time of 2.18 seconds. The XGB achieved 100% accuracy, with a loss of 0.0087 and an overfitting gap of -0.0121%, with a training time of 0.96 seconds. The SVM reached 100% accuracy, with a loss of 0.0704 and an overfitting gap of -0.0342%, with a training time of 4.13 seconds. The MLP achieved 96.67% accuracy, with a loss of 0.2543 and an overfitting gap of -0.0714%, with a training time of 4.28 seconds. Lastly, the LR reached 100% accuracy, with a loss of 0.0539 and an overfitting gap of -0.0438%, with a training time of 4.33 seconds.
@@
Table 2: Performance results of the proposed ALC and other classifiers on the Breast Cancer Wisconsin dataset.
Classifier ALC XGB SVM MLP LR
Training set results
Loss 0.0200 0.0047 0.0388 0.0657 0.0296
Accuracy 0.9890 0.9953 0.9912 0.9890 0.9890
Precision 0.9893 1.0000 0.9912 0.9891 0.9890
Recall 0.9890 1.0000 0.9912 0.9890 0.9890
F1-Score 0.9890 1.0000 0.9912 0.9890 0.9890
Validation set results
Loss 0.0267 0.1164 0.1105 0.0656 0.1067
Accuracy 0.9912 0.8836 0.9825 0.9825 0.9649
Precision 0.9913 0.9561 0.9825 0.9825 0.9658
Recall 0.9912 0.9561 0.9825 0.9825 0.9649
F1-Score 0.9912 0.9560 0.9825 0.9825 0.9651
Overfitting -0.0022% 0.1118% 0.0263% 0.0066% 0.0241%
Time (sec.) 3.27 1.14 3.72 4.65 3.78
Table 2 presents the performance results of the proposed ALC and other classifiers on the Breast Cancer Wisconsin dataset, for both the training and validation sets. Figures 6a and 6b display the loss and accuracy, respectively, on the validation set. The proposed ALC achieved 99.12% accuracy on the validation set, with a loss of 0.0267 and an overfitting gap of -0.0022%, with a training time of 3.27 seconds. The XGB achieved 88.36% accuracy, with a loss of 0.1164 and an overfitting gap of 0.1118%, with a training time of 1.14 seconds. The SVM reached 98.25% accuracy, with a loss of 0.1105 and an overfitting gap of 0.0263%, with a training time of 3.72 seconds. The MLP achieved 98.25% accuracy, with a loss of 0.0656 and an overfitting gap of 0.0066%, with a training time of 4.65 seconds. Lastly, the LR achieved 96.49% accuracy, with a loss of 0.1067 and an overfitting gap of 0.0241%, with a training time of 3.78 seconds.
@@
Table 3: Performance results of the proposed ALC and other classifiers on the Wine dataset.
Classifier ALC XGB SVM MLP LR
Training set results
Loss 0.0000 0.0096 0.0006 0.0783 0.0026
Accuracy 1.0000 0.9904 1.0000 0.9930 1.0000
Precision 1.0000 1.0000 1.0000 0.9931 1.0000
Recall 1.0000 1.0000 1.0000 0.9930 1.0000
F1-Score 1.0000 1.0000 1.0000 0.9930 1.0000
Validation set results
Loss 0.0002 0.0762 0.0001 0.0572 0.0012
Accuracy 1.0000 0.9238 1.0000 1.0000 1.0000
Precision 1.0000 0.9514 1.0000 1.0000 1.0000
Recall 1.0000 0.9444 1.0000 1.0000 1.0000
F1-Score 1.0000 0.9449 1.0000 1.0000 1.0000
Overfitting 0.0000% 0.0666% $0$0070% 0.0000%
Time (sec.) 2.41 1.15 3.94 3.84 3.84
Table 3 presents the performance results of the proposed ALC and other classifiers on the Wine dataset. Figures 7a and 7b display the loss and accuracy, respectively, on the validation set. The proposed ALC achieved 100% accuracy on the validation set, with a loss of 0.0002 and an overfitting gap of 0.0000%, with a training time of 2.41 seconds. The XGB achieved 92.38% accuracy, with a loss of 0.0762 and an overfitting gap of 0.0666%, with a training time of 1.15 seconds. The SVM achieved 100% accuracy, with a loss of 0.0001 and an overfitting gap of 0.0000%, with a training time of 3.94 seconds. The MLP achieved 100% accuracy, with a loss of 0.0572 and an overfitting gap of -0.0070%, with a training time of 3.84 seconds. Lastly, the LR achieved 100% accuracy, with a loss of 0.0012 and an overfitting gap of 0.0000%, with a training time of 3.84 seconds.
@@
Table 4: Performance results of the proposed ALC and other classifiers on the Voice dataset.
Classifier ALC XGB SVM MLP LR
Training set results
Loss 0.0892 0.0012 0.1235 0.0627 0.0928
Accuracy 0.9767 0.9988 0.9775 0.9791 0.9747
Precision 0.9767 1.0000 0.9775 0.9791 0.9748
Recall 0.9767 1.0000 0.9775 0.9791 0.9747
F1-Score 0.9767 1.0000 0.9775 0.9791 0.9747
Validation set results
Loss 0.0677 0.0721 0.1136 0.0618 0.0709
Accuracy 0.9763 0.9279 0.9732 0.9826 0.9811
Precision 0.9766 0.9813 0.9736 0.9827 0.9811
Recall 0.9763 0.9811 0.9732 0.9826 0.9811
F1-Score 0.9764 0.9811 0.9732 0.9827 0.9811
Overfitting 0.0004% 0.0709% 0.0043% -0.0036% -0.0063%
Time (sec.) 3.17 1.24 4.24 12.70 4.44
Table 4 presents the performance results of the proposed ALC and other classifiers on the Voice Gender dataset. Figures 8a and 8b display the log loss and accuracy, respectively, on the validation set. The proposed ALC achieved 97.63% accuracy on the validation set, with a loss of 0.0677 and an overfitting gap of 0.0004%, with a training time of 3.17 seconds. The XGB achieved 92.79% accuracy, with a loss of 0.0721 and an overfitting gap of 0.0709%, with a training time of 1.24 seconds. The SVM achieved 97.32% accuracy, with a loss of 0.1136 and an overfitting gap of 0.0043%, with a training time of 4.24 seconds. The MLP achieved 98.26% accuracy, with a loss of 0.0618 and an overfitting gap of -0.0036%, with a training time of 12.70 seconds. Lastly, the LR achieved 98.11% accuracy, with a loss of 0.0709 and an overfitting gap of -0.0063%, with a training time of 4.44 seconds.
@@
Table 5: Performance results of the proposed ALC and other classifiers on the MNIST dataset.
Classifier ALC XGB SVM MLP LR
Training set results
Loss 0.0000 0.0060 0.0010 0.0404 0.0048
Accuracy 1.0000 0.9940 1.0000 1.0000 1.0000
Precision 1.0000 1.0000 1.0000 1.0000 1.0000
Recall 1.0000 1.0000 1.0000 1.0000 1.0000
F1-Score 1.0000 1.0000 1.0000 1.0000 1.0000
Validation set results
Loss 0.0000 0.0595 0.0082 0.0489 0.0123
Accuracy 0.9975 0.9405 0.9950 0.9900 0.9950
Precision 0.9981 0.9818 0.9953 0.9906 0.9953
Recall 0.9976 0.9800 0.9950 0.9900 0.9950
F1-Score 0.9978 0.9801 0.9950 0.9900 0.9950
Overfitting 0.0025% 0.0535% 0.0050% 0.0100% 0.0050%
Time (sec.) 6.04 2.40 5.46 5.27 5.56
Table 5 presents the performance results of the proposed ALC and other classifiers on the MNIST dataset. Figures 9a and 9b display the log loss and accuracy, respectively, on the validation set. The proposed ALC achieved 99.75% accuracy on the validation set, with a loss of 0.0000 and an overfitting gap of 0.0025%, with a training time of 6.04 seconds. The XGB achieved 94.05% accuracy, with a loss of 0.0595 and an overfitting gap of 0.0535%, with a training time of 2.40 seconds. The SVM achieved 99.50% accuracy, with a loss of 0.0082 and an overfitting gap of 0.0050%, with a training time of 5.46 seconds. The MLP achieved 99.00% accuracy, with a loss of 0.0489 and an overfitting gap of 0.0100%, with a training time of 5.27 seconds. Lastly, the LR achieved 99.50% accuracy, with a loss of 0.0123 and an overfitting gap of 0.0050%, with a training time of 5.56 seconds.

### v2

@@ 5 Results
This section presents the performance results of the proposed ALC on multiple benchmark datasets, as described in Section˜4.1. The experimental parameter settings were configured for each dataset as follows: 500 detoxification cycles, a detoxification power of 10, and dataset-specific numbers of lobules. Specifically, the number of lobules was set to 10 for Iris Flower and Breast Cancer Wisconsin, 15 for Wine and Voice Gender, and 50 for MNIST. The choice of these values was done through a systematic empirical search, where a predetermined set of possible values has been considered, a validation based approach used. Different values of $p$ were searched over each dataset and the value that provided the maximum average classification accuracy on a held-out validation split was selected. This strategy makes sure that the hyperparameters being chosen are tuned in a reproducible and performance-based fashion. In order to have an efficient and fair assessment on the performance of the models, cross-validation was used on the basis of $k$. In particular, each of the datasets has been split into $k$ equal size folds (10-folds were used in our experiments), and the test has been repeated on every fold. Mean results of each of the runs were used to compute final performance metrics. To facilitate later comparison and analysis, additional classifiers, including MLP, SVM, LR, and XGBoost (XGB), were executed on the same datasets. However, all experiments were conducted on an MSI GL63 8RD laptop equipped with an Intel® Core™ i7-8750H × 12 processor and 32 GB of memory. This consistent setup ensured a robust evaluation of the proposed ALC alongside the other classifiers under the same conditions.
@@ 5.1 Performance Metrics
To evaluate the performance of the proposed ALC, several metrics were employed, including log loss (cross-entropy loss), accuracy, precision, recall, F1-score, and training time. Initially, Log loss  Equation˜10 quantifies the divergence between predicted probabilities and actual labels, where lower values indicate better predictive performance [61]. The accuracy  Equation˜11 measures the proportion of correctly classified instances, serving as a straightforward indicator of overall correctness. Moreover, precision  Equation˜12 evaluates the proportion of true positives among all positive predictions, emphasizing the model’s ability to reduce false positives. In contrast, recall  Equation˜13 focuses on the proportion of true positives among all actual positive instances, highlighting the importance of minimizing false negatives. Furthermore, the F1-score  Equation˜14, as the harmonic mean of precision and recall, provides a balanced assessment when class distributions are imbalanced [62]. Moreover, the overfitting gap defined as the difference between training and validation accuracy, provides insights into generalization. A smaller value indicate better generalization, while a larger value indicates overfitting, where the model excels on the training set but struggles with unseen data. Finally, the training time reflects the duration required to train the model, offering insight into its computational efficiency.
$\text{Log Loss}=-\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right)$ (10)
@@
$\text{F1-Score}=2\times\left(\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}\right)$ (14)
@@
5.2 Convergence result of the Training algorithm IFOX
Table˜1 provides the empirical result of the convergence behavior of IFOX based on selected benchmark functions of the CEC2019 suite. Unlike the original FOX, IFOX invariably reaches lower average objective values in most of the test functions, with vastly less variance and greater stability, especially on high-dimensional and multimodal functions like F3 through F7. Furthermore, four recent and competitive optimizers were added to the test in order to prove the effectiveness of IFOX over its predecessor. They consist of GOOSE, ant nesting algorithm (ANA), lagrange elementary optimization (LEO), and FDO, which have shown rather good performance in the literature [63, 64, 65, 66]. The relative performance assures that IFOX attains better convergence characteristics and final solution quality most of the time. Moreover, the gain noticed in performance increase is explained based on adding adaptive inertia control and better search dynamics in IFOX. It is a design that puts an emphasis on more extensive exploration during the initial phases and more targeted exploitation as optimization goes on. Additionally, convergence curves in Figure 4 demonstrate that IFOX converges more rapidly compared to the original FOX and, in addition, obtains better final values. This can be seen by the fact that the fitness trajectory flattened very fast as compared to other methods. Although formal theoretical demonstration of convergence is outside the scope of this paper, the similar results in independent runs (30 runs) give good evidence of the soundness and reliability of IFOX in varied and different optimization landscapes.
Table 1: Comparison of best objective values obtained by the proposed IFOX algorithm across selected CEC2019 benchmark functions over 30 independent runs. The results are compared with those of FOX and several recent optimization algorithms. Lower objective values indicate better performance.
Function/Optimizer F1 F2 F3 F4 F5 F6 F7 F8 F9 F10
GOOSE 1.8E12 6.8E3 13.70 1.60E3 6.09 4.79 274.35 5.57 3.81 20.98
ANA - 4.00 13.70 38.50 1.22 - 116.59 5.47 2.00 2.71
LEO 7.3E9 17.47 12.70 69.86 2.76 3.01 195.56 5.06 3.26 20.01
FDO 4585.27 4.00 13.7 34.08 2.13 12.13 120.40 6.1 2.01 2.71
FOX 1.00 4.72 9.88 147.21 5.13 298.10 1.017 1.38 1.41 21.49
IFOX 1.00 4.61 2.42 35.80 1.88 1.00 1.00 1.24 1.33 20.99
In order to compare performance on the whole, each optimization algorithm was ranked by the number of functions with the best objective value. The overall and the average ranking of all functions is given in Table˜2. IFOX presented a minimum total (18) and average (1.8) rank and maintained a better performance as compared to the rest of the optimization algorithms. Due to IFOX’s superior convergence and stability, it was chosen as the training algorithm in this study.
Table 2: Function-wise ranks, total rank, and average rank for each optimization algorithm across the 10 selected CEC2019 benchmark functions. Lower ranks indicate better performance.
Function/Optimizer F1 F2 F3 F4 F5 F6 F7 F8 F9 F10 Total Rank Avg. Rank
IFOX 1 3 1 2 2 1 1 1 1 5 18 1.8
FDO 3 1 4 1 3 4 4 6 4 1 31 3.1
ANA 6 1 4 3 1 6 3 4 3 1 32 3.2
FOX 1 4 2 5 5 5 2 2 2 6 34 3.4
LEO 4 5 3 4 4 2 5 3 5 3 38 3.8
GOOSE 6 6 4 6 6 3 6 5 6 4 52 5.2
5.3 Experimental results of ALC
The performance results of the proposed ALC are presents through this subsection, summarized in the figures and tables. Additionally, comparisons with other classifiers, including MLP, SVM, LR, and XGB, have been conducted on the five datasets described in Section˜4.1.
Table 3 presents the performance results of the proposed ALC and other classifiers on the Iris Flower dataset. Additionally, Figures 5(a) and 5(b) show the loss and accuracy, respectively, across the validation folds. The proposed ALC achieved 100% accuracy with a loss of 0.0169, an overfitting gap of $0.0231\%$, and a training time of 2.12 seconds. The XGB also achieved 100% accuracy with a loss of 0.0085, an overfitting gap of $0.0144\%$, and a training time of 0.91 seconds. Similarly, the SVM reached 100% accuracy with a loss of 0.0704, an overfitting gap of $0.0384\%$, and a training time of 4.42 seconds. The MLP attained 96.67% accuracy with a loss of 0.2417, an overfitting gap of $0.0714\%$, and a training time of 4.18 seconds. Lastly, the LR reached 100% accuracy with a loss of 0.0543, an overfitting gap of $0.0415\%$, and a training time of 4.31 seconds.
Table 3: Cross-validation performance of the proposed ALC and other classifiers on the Iris Flower dataset (mean over 10-folds).
Metric ALC XGB SVM MLP LR
Loss 0.0169 0.0085 0.0691 0.2417 0.0543
@@ 5.3 Experimental results of ALC
Overfitting -0.0231% -0.0144% -0.0384% -0.0689% -0.0415%
Time (sec.) 2.12 0.91 4.42 4.18 4.31
(a) Log loss on the Iris dataset
(b) Accuracy on the Iris dataset
Table 4 presents the performance results of the proposed ALC and other classifiers on the Breast Cancer Wisconsin dataset. Figures 6(a) and 6(b) display the loss and accuracy, respectively, on the validation folds. The proposed ALC achieved 99.12% accuracy, with a loss of 0.0261 and an overfitting gap of -0.0029%, with a training time of 3.62 seconds. The XGB achieved 88.36% accuracy, with a loss of 0.1132 and an overfitting gap of 0.1178%, with a training time of 1.09 seconds. The SVM reached 98.25% accuracy, with a loss of 0.1105 and an overfitting gap of 0.0213%, with a training time of 3.79 seconds. The MLP achieved 98.25% accuracy, with a loss of 0.0682 and an overfitting gap of 0.0086%, with a training time of 4.53 seconds. Lastly, the LR achieved 96.49% accuracy, with a loss of 0.1319 and an overfitting gap of 0.0237%, with a training time of 3.81 seconds.
Table 4: Cross-validation performance of the proposed ALC and other classifiers on the Breast Cancer Wisconsin dataset (mean over 10-folds).
Metric ALC XGB SVM MLP LR
Loss 0.0261 0.1132 0.1203 0.0682 0.1319
Accuracy 0.9932 0.8927 0.9833 0.9833 0.9559
Precision 0.9943 0.9233 0.9833 0.9833 0.9638
Recall 0.9932 0.9233 0.9833 0.9833 0.9559
F1-Score 0.9932 0.9260 0.9833 0.9833 0.9631
Overfitting -0.0029% 0.1178% 0.0213% 0.0086% 0.0237%
Time (sec.) 3.62 1.09 3.79 4.53 3.81
@@
Table 5 presents the performance results of the proposed ALC and other classifiers on the Wine dataset. Figures 7(a) and 7(b) display the loss and accuracy, respectively, on the validation folds. The proposed ALC achieved 100% accuracy, with a loss of 0.0011 and an overfitting gap of 0.0000%, with a training time of 2.38 seconds. The XGB achieved 92.38% accuracy, with a loss of 0.0691 and an overfitting gap of 0.0675%, with a training time of 1.17 seconds. The SVM achieved 100% accuracy, with a loss of 0.0001 and an overfitting gap of 0.0000%, with a training time of 3.89 seconds. The MLP achieved 100% accuracy, with a loss of 0.0568 and an overfitting gap of -0.0068%, with a training time of 3.89 seconds. Lastly, the LR achieved 100% accuracy, with a loss of 0.0012 and an overfitting gap of 0.0000%, with a training time of 3.92 seconds.
Table 5: Cross-validation performance of the proposed ALC and other classifiers on the Wine dataset (mean over 10-folds).
Metric ALC XGB SVM MLP LR
Loss 0.0011 0.0691 0.0001 0.0568 0.0012
Accuracy 1.0000 0.9258 1.0000 1.0000 1.0000
Precision 1.0000 0.9534 1.0000 1.0000 1.0000
Recall 1.0000 0.9424 1.0000 1.0000 1.0000
F1-Score 1.0000 0.9429 1.0000 1.0000 1.0000
Overfitting 0.0000% 0.0675% 0.0000% -0.0068% 0.0000%
Time (sec.) 2.38 1.17 3.89 3.89 3.92
@@
Table 6 presents the performance results of the proposed ALC and other classifiers on the Voice Gender dataset. Figures 8(a) and 8(b) display the log loss and accuracy, respectively, on the validation folds. The proposed ALC achieved 97.63% accuracy, with a loss of 0.0613 and an overfitting gap of 0.0004%, with a training time of 3.21 seconds. The XGB achieved 92.79% accuracy, with a loss of 0.0706 and an overfitting gap of 0.0601%, with a training time of 1.19 seconds. The SVM achieved 97.32% accuracy, with a loss of 0.1932 and an overfitting gap of 0.0043%, with a training time of 4.35 seconds. The MLP achieved 98.26% accuracy, with a loss of 0.0622 and an overfitting gap of -0.0036%, with a training time of 12.59 seconds. Lastly, the LR achieved 98.11% accuracy, with a loss of 0.0601 and an overfitting gap of -0.0063%, with a training time of 4.61 seconds.
Table 6: Cross-validation performance of the proposed ALC and other classifiers on the Voice dataset (mean over 10-folds).
Metric ALC XGB SVM MLP LR
Loss 0.0613 0.0706 0.1932 0.0622 0.0601
Accuracy 0.9752 0.9164 0.9750 0.9800 0.9811
Precision 0.9750 0.9325 0.9736 0.9811 0.9811
Recall 0.9752 0.9105 0.9750 0.9800 0.9811
F1-Score 0.9752 0.9150 0.9750 0.9811 0.9811
Overfitting 0.0004% 0.0601% 0.0043% -0.0036% -0.0063%
Time (sec.) 3.21 1.19 4.35 12.59 4.61
@@
Table 7 presents the performance results of the proposed ALC and other classifiers on the MNIST dataset. Figures 9(a) and 9(b) display the log loss and accuracy, respectively, on the validation set. The proposed ALC achieved 99.75% accuracy on the validation set, with a loss of 0.0000 and an overfitting gap of 0.0025%, with a training time of 6.18 seconds. The XGB achieved 94.05% accuracy, with a loss of 0.0581 and an overfitting gap of 0.0571%, with a training time of 2.35 seconds. The SVM achieved 99.50% accuracy, with a loss of 0.0076 and an overfitting gap of 0.0050%, with a training time of 5.38 seconds. The MLP achieved 99.00% accuracy, with a loss of 0.0473 and an overfitting gap of 0.0100%, with a training time of 5.22 seconds. Lastly, the LR achieved 99.50% accuracy, with a loss of 0.0137 and an overfitting gap of 0.0050%, with a training time of 5.61 seconds.
Table 7: Cross-validation performance of the proposed ALC and other classifiers on the MNIST dataset (mean over 10-folds).
Metric ALC XGB SVM MLP LR
Loss 0.0000 0.0581 0.0076 0.0473 0.0137
Accuracy 0.9975 0.9421 0.9967 0.9900 0.9967
Precision 0.9970 0.9828 0.9953 0.9906 0.9953
Recall 0.9967 0.9802 0.9967 0.9900 0.9967
F1-Score 0.9987 0.9800 0.9967 0.9900 0.9967
Overfitting 0.0025% 0.0571% 0.0050% 0.0100% 0.0050%
Time (sec.) 6.18 2.35 5.38 5.22 5.61